# PETLP: A Privacy-by-Design Pipeline for Social Media Data in AI Research

### v1

@@ Abstract
Social media data presents AI researchers with overlapping obligations under the GDPR, copyright law, and platform terms – yet existing frameworks fail to integrate these regulatory domains, leaving researchers without unified guidance. We introduce PETLP (Privacy-by-design Extract, Transform, Load, and Present), a compliance framework that embeds legal safeguards directly into extended ETL pipelines. Central to PETLP is treating Data Protection Impact Assessments as living documents that evolve from pre-registration through dissemination. Through systematic Reddit analysis, we demonstrate how extraction rights fundamentally differ between qualifying research organisations (who can invoke DSM Article 3 to override platform restrictions) and commercial entities (bound by terms of service), whilst GDPR obligations apply universally. We reveal why true anonymisation remains unachievable for social media data and expose the legal gap between permitted dataset creation and uncertain model distribution. By structuring compliance decisions into practical workflows and simplifying institutional data management plans, PETLP enables researchers to navigate regulatory complexity with confidence, bridging the gap between legal requirements and research practice.
@@ 1 Introduction
Social media platforms have become essential data sources for computational and social science research, enabling investigations into political movements (Solovev and Pröllochs 2022; Iqbal et al. 2022), labour inequalities (Xu 2024), discrimination (Nesterov, Hollink, and van Ossenbruggen 2024) and cultural phenomena (Nguyen et al. 2023). These platforms provide datasets of unprecedented scale and immediacy, offering insights into societal trends and human behaviour that surpass traditional methods (Bundtzen 2023). Concurrently, methodological advances – from BERT-based models (Ji et al. 2020; Ananthakrishnan et al. 2022) to Large Language Models (LLMs) (Deng et al. 2023; Alhamed, Ive, and Specia 2024; Vuruma et al. 2024) – have expanded analytical possibilities whilst intensifying ethical and regulatory challenges.
@@
The research landscape presents AI researchers with multiple layers of complexity. At the foundational level, the definition of social media itself continues to evolve – conventional websites now incorporate social features like user profiles, messaging, and content-sharing to enhance engagement (Garcia-Pueyo et al. 2023; He et al. 2024), blurring boundaries and creating ambiguity about applicable legal and ethical frameworks (Terzis, Veale, and Gaumann 2024). Regulatory frameworks add a second layer of uncertainty: the GDPR offers limited guidance on platform-to-researcher data sharing (European Digital Media Observatory 2022) whilst employing broad, often undefined concepts such as ‘personal data’ and ‘scientific research’.
@@
These definitional and regulatory ambiguities intersect with a third layer – platform governance tensions. Platforms increasingly invoke privacy concerns to restrict data access (Bruns 2021; Morten, Nicholas, and Viljoen 2024), despite legal precedents establishing no reasonable expectation of privacy for publicly accessible content (Rom 2010). Nevertheless, the ethical landscape presents equally significant challenges. Empirical studies demonstrate that fewer than 10% of social media research publications address ethical implications beyond securing basic Institutional Review Board (IRB) approval (Taylor and Pagliari 2018; Chiauzzi and Wicks 2019). Together, these layers create a reinforcing cycle: definitional uncertainty enables regulatory gaps, which justify platform restrictions, ultimately exposing ethical inadequacies that undermine public trust and research value (European Digital Media Observatory 2022).
@@
To navigate these challenges, we propose the Privacy-by-design Extract, Transform, Load and Present (PETLP) framework. Unlike conventional data pipelines that treat compliance as an afterthought, PETLP directly addresses each layer of complexity: it clarifies definitional ambiguities through structured extraction typologies, operationalises vague regulatory concepts via living DPIAs, navigates platform restrictions through legally-grounded alternatives, and elevates ethical practice beyond checkbox compliance. The framework integrates three intersecting legal domains – data protection regulations, intellectual property (IP) rights, and platform contract law – transforming regulatory requirements from external constraints into structured technical workflows. PETLP addresses the fundamental question: How can researchers responsibly access, process, and present social media data and derivative models whilst fulfilling legal obligations throughout the data pipeline?
@@ 2 Background and Scope
Three intersecting legal frameworks – data protection, intellectual property, and contract law – collectively govern how AI researchers may access, process, and use social media data. Understanding these frameworks is essential, as they shape every decision from initial data collection through model deployment. This section establishes the conceptual and legal foundations necessary for responsible AI research, first defining social media data and its distinctive characteristics that create unique research challenges, then examining each legal framework in detail.
Our analysis centres on European Union law, which provides a comprehensive, principle-based model of data governance (Lynskey 2015; European Commission 2021). The EU offers established legal regimes with mature interpretive guidance, making it an instructive model for researchers globally. While we focus on current frameworks, researchers should note that the EU AI Act (effective August 2025) influences development practices – Risk Management Systems (Article 9), Data Governance (Article 10), Technical Documentation (Article 11), and General Purpose AI provider obligations (Articles 50 and 53) reinforce principles of transparency, documentation, and risk management central to responsible research. Appendix C explores potential AI Act integration with PETLP. Researchers operating under other jurisdictions should consult Brown et al. (2024) for US-specific web scraping guidance and AGORA (Arnold et al. 2024) for cross-jurisdictional AI governance instruments.
@@ 2.1 Defining Social Media Data
Social media data refers to content generated and collected from users and their interactions within platforms, often in the context of social relationships [^1] The European Data Protection Board (EDPB) defines social media as ’online platforms that enable the development of networks and communities of users, among which information and content is shared’, where key characteristics include account creation, content sharing, and user connections (European Data Protection Supervisor 2021, p. 4). (Olteanu et al. 2019; Stieglitz et al. 2018; Batrinca and Treleaven 2015). This encompasses traditional social networking sites like Facebook (now Meta) and Twitter (now X), alongside content-sharing platforms (YouTube, Instagram), messaging applications with social features (WhatsApp, WeChat), and community-based platforms (Reddit, Discord) (Boyd and Ellison 2007; Obar and Wildman 2015).
This social dimension creates unique legal and ethical challenges, as individuals can be identified not merely through their own posts but through their networks of connections, interaction patterns, and community memberships – even when traditional identifiers are removed. A user’s participation in specific communities, commenting patterns, or social graph can serve as digital fingerprints, making anonymisation particularly difficult and raising the stakes for privacy protection.
@@ 2.2 Legal Foundations for Social Media Data Pipelines
Navigating social media data for AI training requires researchers to traverse a landscape marked by fundamental legal uncertainty (Longpre et al. 2024; Dornis and Stober 2025). Three distinct yet interconnected legal regimes govern this terrain, each imposing specific obligations that shape every stage of the research pipeline. Understanding these foundations is essential – not merely for compliance, but for designing research that can withstand scrutiny, maintain access to data sources, and preserve public trust.
The GDPR establishes comprehensive privacy principles yet provides limited operational guidance for research contexts. Key concepts remain broadly defined: what constitutes ‘research purposes’, how ‘personal data’ extends in social media contexts, or which activities qualify as ‘scientific research’ (European Digital Media Observatory 2022). Similarly, intellectual property law presents ambiguity – whilst the EU’s Text and Data Mining (TDM) exceptions permit researchers to extract patterns from protected data, it remains unclear whether these protections extend to training AI models. Contract law adds further complexity through platform Terms of Service (ToS) that often prohibit automated access, redistribution, or AI training, creating potential conflicts with statutory research rights.
These overlapping regimes create compound risks. A single research project may simultaneously navigate data subject rights under the GDPR, copyright protections on user-generated content, database rights held by platforms, and contractual restrictions imposed through ToS. The following sections examine each legal domain in detail, establishing the conceptual foundations necessary for understanding their technical implications and practical intersections throughout the research pipeline.
2.2.1 GENERAL DATA PROTECTION REGULATION
@@
Three overarching principles guide GDPR obligations. First, necessity and proportionality require that data processing be both essential for research objectives and balanced against privacy impacts. Researchers cannot scrape entire subreddits just in case some posts prove useful – each data point must serve the stated research purpose. Second, accountability demands that researchers not only comply with the GDPR but demonstrate compliance through documentation and safeguards. What constitutes ‘appropriate’ safeguards depends on risk: analysing public product reviews requires basic security measures, while processing mental health discussions demands careful handling. Third, rights and freedoms of data subjects remain paramount – individuals retain rights to access, correct, or object to processing of their data, though Article 89 allows some limitations for scientific research where these rights would ‘seriously impair’ research objectives. Article 5 translates these into six specific principles, while Article 25 mandates that these protections be embedded into research design from the outset – privacy by design, not by afterthought.
@@
Social media data fundamentally qualifies as personal data under the GDPR, irrespective of public accessibility. The determining factor lies not in the data’s availability but in its potential to identify individuals directly or indirectly (Article 29 Data Protection Working Party 2017; Court of Justice of the European Union 2024b). Under the GDPR, data is rendered anonymous only when individuals cannot be identified by any means ‘reasonably likely to be used’ (Recital 26), accounting for future technological advances and data linkage possibilities – a standard rarely achievable with social media’s rich behavioural patterns. Moreover, social media data frequently enables inference of special category data under Article 9; recent Court of Justice of the European Union (CJEU) rulings confirm that the mere possibility of inferring protected attributes – regardless of accuracy – activates heightened safeguards (CJEU 2022, 2023b, 2023a). This is particularly significant for AI research, where models may inadvertently memorise training data or develop unforeseen inference capabilities (European Data Protection Board 2024b), demanding researchers account for both intended outcomes and latent model capacities.
The Researcher as Joint Controller
@@
While research institutions formally act as controllers, controllership is functionally determined by actual influence over data processing decisions (European Data Protection Board 2020). AI researchers directly determine collection methodologies, transformation techniques, and model training approaches – decisions integral to defining the means of processing under Article 4(7). This paper therefore advocates viewing researchers as joint controllers with their institutions, recognising that joint controllership emerges when parties jointly determine purposes and means through complementary decisions (European Data Protection Board 2020). This perspective ensures privacy considerations permeate the entire research pipeline rather than remaining abstract institutional obligations, creating accountability where data handling decisions are actually made (Appendix D, Figure 2).
Data Protection by Design and Impact Assessments
@@
For AI researchers using social media data, DPIAs are effectively mandatory under Article 35, as such research routinely triggers multiple high-risk criteria identified by the Article 29 Data Protection Working Party (2017): large-scale processing, dataset combination, and innovative technology deployment (Appendix B). A compliant DPIA must: (1) systematically describe the processing including data flows and infrastructure (Article 35(7)(a)); (2) assess necessity and proportionality while demonstrating Article 5 principle implementation (35(7)(b)); and (3) identify risks with appropriate mitigations (35(7)(c)). Notably, DPIAs are living documents requiring updates throughout iterative AI development cycles (Article 29 Data Protection Working Party 2017, p. 9) (Figure 1).
Establishing a Legal Basis for Processing
@@
Social media researchers typically rely on either public interest (Article 6(1)(e)) or legitimate interests (Article 6(1)(f)) as their legal basis, with the choice determined by organisational context (Appendix D, Figure 3). Public authorities and universities operating under statutory research mandates can invoke public interest grounds, provided they demonstrate that their research serves a recognised public benefit (Court of Justice of the European Union 2024a). In contrast, private sector entities must conduct and document a Legitimate Interest Assessment (LIA) to justify their processing activities, as outlined in the Article 29 Data Protection Working Party (2017). Both require distinguishing broader research interests from specific processing purposes. The GDPR affords research significant flexibilities under Article 89 – including purpose compatibility, extended retention, and conditional rights limitations – provided researchers implement appropriate safeguards (European Data Protection Board 2021). These privileges enable research while maintaining privacy protection through proportionate technical and organisational measures.
2.2.2 INTELLECTUAL PROPERTY AND CONTRACT LAW
@@
Copyright and Database Rights
Under the InfoSoc Directive (Directorate-General for Communications Networks, Content and Technology 2001), copyright grants authors exclusive economic rights over reproduction, distribution and public communication of their works. The CJEU has set a low originality threshold – ‘the author’s own intellectual creation’ – holding in Infopaq (C-5/08) that even eleven-word extracts could qualify if bearing the author’s personal stamp, and in Painer (C-145/10) that portrait photographs merit protection through creative choices in framing and lighting. This means even modest user-generated content (posts, short videos) may be protected if demonstrating individualised creative input. Simultaneously, the Database Directive (Council of the European Union 1996) establishes sui generis rights where ‘substantial investment’ exists in obtaining, verifying or presenting contents, potentially protecting platforms’ aggregated collections of posts and metadata. This dual-layer protection – individual posts via copyright, aggregated collections via database rights – creates complex questions about scraping and reusing social media datasets, particularly where data provenance is unclear (OECD 2025). To balance these rights with scientific advancement, the DSM Directive (Council of the European Union 2019) introduced mandatory Text and Data Mining (TDM) exceptions: Article 3 grants research organisations unwaivable TDM rights for scientific research on lawfully accessed content, whilst Article 4 provides broader TDM rights that rightholders may reserve through machine-readable means such as the Robot Exclusion Protocol (robots.txt) – a text file that specifies which parts of a website automated crawlers may access.
Contractual and Technical Safeguards
Contract law introduces an additional layer of legal obligations, often in the form of platform ToS or API agreements. These contractual terms define the conditions under which users, including researchers, may access and interact with platform data. Violating these terms can expose researchers to contractual liability, including account suspension, loss of data access, or even legal action for breach of contract. Whilst the DSM Directive renders certain contractual restrictions on research unenforceable, many limitations remain valid and binding. Platforms also implement robots.txt to control automated access, which researchers should respect as both a legal requirement (where they constitute Article 4 DSM reservations) and an ethical consideration.

### v2

@@ Abstract
Social media data presents AI researchers with overlapping obligations under the GDPR, copyright law, and platform terms – yet existing frameworks fail to integrate these regulatory domains, leaving researchers without unified guidance. We introduce PETLP (Privacy-by-design Extract, Transform, Load, and Present), a compliance framework that embeds legal safeguards directly into extended ETL pipelines. Central to PETLP is treating Data Protection Impact Assessments as living documents that evolve from pre-registration through dissemination. Through systematic Reddit analysis, we demonstrate how extraction rights fundamentally differ between qualifying research organisations (who can invoke DSM Article 3 to override platform restrictions) and commercial entities (bound by terms of service), whilst GDPR obligations apply universally. We demonstrate why true anonymisation remains unachievable for social media data and expose the legal gap between permitted dataset creation and uncertain model distribution. By structuring compliance decisions into practical workflows and simplifying institutional data management plans, PETLP enables researchers to navigate regulatory complexity with confidence, bridging the gap between legal requirements and research practice.
@@ 1 Introduction
Social media platforms have become essential data sources for computational and social science research, enabling investigations into political movements (Solovev and Pröllochs 2022; Iqbal et al. 2022), labour inequalities (Xu 2024), discrimination (Nesterov, Hollink, and van Ossenbruggen 2024) and cultural phenomena (Nguyen et al. 2023). These platforms provide datasets of unprecedented scale and immediacy, offering insights into societal trends and human behaviour that surpass traditional methods (Bundtzen 2023). Concurrently, methodological advances – from BERT-based models (Ji et al. 2020; Ananthakrishnan et al. 2022) to Large Language Models (LLMs) (Deng et al. 2023; Alhamed, Ive, and Specia 2024; Vuruma et al. 2024) – have expanded analytical possibilities, yet with a cost of intensifying ethical and regulatory challenges.
@@
The research landscape presents AI researchers with multiple layers of complexity. At the foundational level, the definition of social media itself continues to evolve – conventional websites now incorporate social features like user profiles, messaging, and content-sharing to enhance engagement (Garcia-Pueyo et al. 2023; He et al. 2024), blurring boundaries and creating ambiguity about applicable legal and ethical frameworks (Terzis, Veale, and Gaumann 2024). Regulatory frameworks add a second layer of uncertainty: the GDPR offers limited guidance on platform-to-researcher data sharing (European Digital Media Observatory 2022) while employing broad, often undefined concepts such as ‘personal data’ and ‘scientific research’.
@@
These definitional and regulatory ambiguities intersect with a third layer – platform governance tensions. Platforms increasingly invoke privacy concerns to restrict data access (Bruns 2021; Morten, Nicholas, and Viljoen 2024), despite legal precedents establishing no reasonable expectation of privacy for publicly accessible content (Rom 2010). Nevertheless, the ethical landscape presents equally significant challenges, with empirical studies demonstrating that fewer than 10% of social media research publications address ethical implications beyond securing basic Institutional Review Board (IRB) approval (Taylor and Pagliari 2018; Chiauzzi and Wicks 2019). The intersection of these three layers – evolving platform definitions, regulatory gaps, and increasingly restrictive platform policies – creates a fragmented compliance landscape where researchers face contradictory obligations that ultimately compromise both research quality and public confidence (European Digital Media Observatory 2022).
@@
To navigate these challenges, we propose the Privacy-by-design Extract, Transform, Load and Present (PETLP) framework. Unlike conventional data pipelines that treat compliance as an afterthought, PETLP directly addresses each layer of complexity: it clarifies definitional ambiguities, operationalises vague regulatory concepts via living DPIAs, navigates platform restrictions through legally-grounded alternatives, and elevates ethical practice beyond checkbox compliance. The framework integrates three intersecting legal domains – data protection regulations, intellectual property (IP) rights, and platform contract law – transforming regulatory requirements from external constraints into structured technical workflows. PETLP addresses the fundamental question: How can researchers responsibly access, process, and present social media data and derivative models whilst fulfilling legal obligations throughout the data pipeline?
@@ 2 Background and Scope
Social media data presents unique challenges for AI research. Its distinctive characteristics, including mixed public-private boundaries, embedded personal information, and complex ownership structures, necessitate careful legal consideration before collection, processing, or model deployment. Understanding the legal frameworks governing such data is therefore essential for responsible AI research practice.
Our analysis examines three intersecting legal domains that collectively govern AI researchers’ use of social media data: data protection, IP, and contract law. We centre this analysis on European Union law, which provides a comprehensive, principle-based governance model (Lynskey 2015; European Commission 2021). The GDPR’s extraterritorial reach – applying to any processing of EU-located individuals’ data regardless of researchers’ location – makes compliance virtually certain for social media AI research, as global platforms inevitably include EU users’ content. This broad applicability establishes EU law as the de facto baseline for international research standards.
Whilst focusing on current requirements, we acknowledge the EU AI Act (effective August 2025) introduces complementary development-phase obligations that reinforce transparency and risk management principles (see Appendix C). Researchers outside Europe should consult jurisdiction-specific guidance, such as Brown et al. (2024) for US web scraping or the AI Governance and Regulatory Archive (AGORA) (Arnold et al. 2024) for cross-jurisdictional analysis. However, GDPR’s extraterritorial application means most social media AI research requires European compliance regardless of researcher location, making our framework broadly applicable.
@@ 2.1 Defining Social Media Data
Social media data refers to content generated and collected from users and their interactions within platforms, often in the context of social relationships [^1] The European Data Protection Board (EDPB) defines social media as ‘online platforms that enable the development of networks and communities of users, among which information and content is shared’, where key characteristics include account creation, content sharing, and user connections (European Data Protection Supervisor 2021, p. 4). (Olteanu et al. 2019; Stieglitz et al. 2018; Batrinca and Treleaven 2015). This encompasses traditional social networking sites like Facebook (now Meta) and Twitter (now X), alongside content-sharing platforms (YouTube, Instagram), messaging applications with social features (WhatsApp, WeChat), and community-based platforms (Reddit, Discord) (Boyd and Ellison 2007; Obar and Wildman 2015).
This social dimension creates unique legal and ethical challenges, as individuals can be identified not merely through their own posts but through their networks of connections, interaction patterns, and community memberships – even when traditional identifiers are removed. A user’s participation in specific communities, commenting patterns, or social graph can serve as digital fingerprints, making anonymisation particularly difficult and raising the stakes for privacy protection (Qian et al. 2019; Jiang et al. 2018; Tian et al. 2018).
@@ 2.2 Legal Foundations for Social Media Data Pipelines
Utilising social media data for AI training requires researchers to traverse a landscape marked by fundamental legal uncertainty (Longpre et al. 2024; Dornis and Stober 2025). Three distinct yet interconnected legal regimes govern this terrain, each imposing specific obligations that shape every stage of the research pipeline. Understanding these foundations is essential – not merely for compliance, but for designing research that can withstand scrutiny, maintain access to data sources, and preserve public trust.
The GDPR establishes comprehensive privacy principles yet provides limited operational guidance for research contexts. Key concepts remain broadly defined (European Digital Media Observatory 2022): what constitutes ‘research purposes’, how ‘personal data’ extends in social media contexts, or which activities qualify as ‘scientific research’. Similarly, IP law presents ambiguity – whilst the EU’s Text and Data Mining (TDM) exceptions permit researchers to extract patterns from copyright protected data, it remains unclear whether these protections extend to training AI models. Contract law adds further complexity through platform Terms of Service (ToS) that often prohibit automated access, redistribution, or AI training, creating potential conflicts with statutory research rights.
These overlapping regimes create compound risks. A single research project will typically navigate data subject rights under the GDPR, copyright protections on user-generated content, database rights held by platforms, and contractual restrictions imposed through ToS. The following sections examine each legal domain in detail, establishing the conceptual foundations necessary for understanding their technical implications and practical intersections throughout the research pipeline.
GDPR
@@
Three overarching principles guide GDPR obligations. First, necessity and proportionality require that data processing be both essential for research objectives and balanced against privacy impacts. Researchers cannot scrape entire subreddits just in case some posts prove useful – each data point must serve the stated research purpose. Second, accountability demands that researchers not only comply with the GDPR but demonstrate compliance through documentation and safeguards. What constitutes ‘appropriate’ safeguards depends on risk: analysing public product reviews requires basic security measures, while processing mental health discussions demands careful handling. Third, rights and freedoms of data subjects remain paramount – individuals retain rights to access, correct, or object to processing of their data, though Article 89 allows some limitations for scientific research where these rights would ‘seriously impair’ research objectives. Article 5 translates these into six specific principles, and Article 25 mandates that these protections be embedded into research design from the outset – privacy by design, not by afterthought.
@@
Social Media Data as Personal Data Social media data qualifies as personal data under the GDPR, irrespective of public accessibility. The determining factor lies not in the data’s availability but in its potential to identify individuals directly or indirectly (Article 29 Data Protection Working Party 2017; Court of Justice of the European Union 2024b). Under the GDPR, data is rendered anonymous only when individuals cannot be identified by any means ‘reasonably likely to be used’ (Recital 26), accounting for future technological advances and data linkage possibilities – a standard rarely achievable with social media’s rich behavioural patterns. Moreover, social media data frequently enables inference of special category data under Article 9; recent Court of Justice of the European Union (CJEU) rulings confirm that the mere possibility of inferring protected attributes activates heightened safeguards regardless of accuracy (Court of Justice of the European Union 2022, 2023b, 2023a). This is particularly significant for AI research, where models may inadvertently memorise training data or develop unforeseen inference capabilities (European Data Protection Board 2024b), demanding researchers account for both intended outcomes and latent model capacities.
@@
The Researcher as Joint Controller While research institutions formally act as controllers, controllership is functionally determined by actual influence over data processing decisions (European Data Protection Board 2020). AI researchers directly determine collection methodologies, transformation techniques, and model training approaches – decisions integral to defining the means of processing under Article 4(7). This paper therefore advocates viewing researchers as joint controllers with their institutions, recognising that joint controllership emerges when parties jointly determine purposes and means through complementary decisions (European Data Protection Board 2020). This perspective ensures privacy considerations permeate the entire research pipeline rather than remaining abstract institutional obligations, creating accountability where data handling decisions are actually made (Appendix D, Figure 2).
@@
Data Protection by Design and Impact Assessments For AI researchers using social media data, DPIAs are effectively mandatory under Article 35, as such research routinely triggers multiple high-risk criteria identified by the Article 29 Data Protection Working Party (2017): large-scale processing, dataset combination, and innovative technology deployment (Appendix B). A compliant DPIA must: (1) systematically describe the processing including data flows and infrastructure (Article 35(7)(a)); (2) assess necessity and proportionality while demonstrating Article 5 principle implementation (35(7)(b)); and (3) identify risks with appropriate mitigations (35(7)(c)). Notably, DPIAs are living documents requiring updates throughout iterative AI development cycles (Article 29 Data Protection Working Party 2017, p. 9).
@@
Establishing a Legal Basis for Processing Social media researchers typically rely on either public interest (Article 6(1)(e)) or legitimate interests (Article 6(1)(f)) as their legal basis, with the choice determined by organisational context (Appendix D, Figure 3). Public authorities and universities operating under statutory research mandates can invoke public interest grounds, provided they demonstrate that their research serves a recognised public benefit (Court of Justice of the European Union 2024a). In contrast, private sector entities must conduct and document a Legitimate Interest Assessment (LIA) to justify their processing activities, as outlined in the Article 29 Data Protection Working Party (2017). Both require distinguishing broader research interests from specific processing purposes. The GDPR affords research significant flexibilities under Article 89 – including purpose compatibility, extended retention, and conditional rights limitations – provided researchers implement appropriate safeguards (European Data Protection Board 2021). These privileges enable research while maintaining privacy protection through proportionate technical and organisational measures.
IP and Contract Law
@@
Copyright and Database Rights Under the InfoSoc Directive (Directorate-General for Communications Networks, Content and Technology 2001), copyright grants authors exclusive economic rights over reproduction, distribution and public communication of their works. The CJEU has set a low originality threshold – ‘the author’s own intellectual creation’ – holding in Infopaq International A/S v Danske Dagblades Forening (C-5/08) that even eleven-word extracts could qualify if bearing the author’s personal stamp, and in Painer v Standard Verlags GmbH (C-145/10) that portrait photographs merit protection through creative choices in framing and lighting. This means even modest user-generated content (posts, short videos) may be protected if demonstrating individualised creative input. Simultaneously, the Database Directive (Council of the European Union 1996) establishes sui generis rights where ‘substantial investment’ exists in obtaining, verifying or presenting contents, potentially protecting platforms’ aggregated collections of posts and metadata. This dual-layer protection – individual posts via copyright, aggregated collections via database rights – creates complex questions about scraping and reusing social media datasets, particularly where data provenance is unclear (OECD 2025). To balance these rights with scientific advancement, the DSM Directive (Council of the European Union 2019) introduced mandatory Text and Data Mining (TDM) exceptions: Article 3 grants research organisations unwaivable TDM rights for scientific research on lawfully accessed content, while Article 4 provides broader TDM rights that rightholders may reserve through machine-readable means such as the Robot Exclusion Protocol (robots.txt) – a text file that specifies which parts of a website automated crawlers may access.
Contractual and Technical Safeguards Contract law introduces an additional layer of legal obligations, often in the form of platform ToS or API agreements. These contractual terms define the conditions under which users, including researchers, may access and interact with platform data. Violating these terms can expose researchers to contractual liability, including account suspension, loss of data access, or even legal action for breach of contract. Whilst the DSM Directive renders certain contractual restrictions on research unenforceable, many limitations remain valid and binding. Platforms also implement robots.txt to control automated access, which researchers should respect as both a legal requirement (where they constitute Article 4 DSM reservations) and an ethical consideration.